the taylor family has a turnout of 38 people. they decidedd to go ice skating instead of having a picnic. admission for the group cost $134.00. How many adults and how many children were at the reunion?
prices
adults 4.00
kids 3.00

Answer by stanbon(75887)

Let number of adults be "x"
The number of kids is "38-x"
Cost of adult tickets is 4x
Cost of kids tickets is 3(38-x)
EQUATION:
Cost + Cost = $134
4x+3(38-x)=134
4x+114-3x=134
x=20 (number of adults)
38-x=18 (number of Kids)
Cheers,
Stan H.